NNT Ltd pays no dividends and has a current stock price of $20 per share. Its returns have an annual volatility of 15% and the risk free interest rate is 4% per annum. (i). Calculate the value of a one-year at-the-money call option using the Black Scholes model.
Answer #1

d1=(ln(S/X)+(risk free rate+volatility^2/2)*t)/(volatility*sqrt(t))=(4%+(15%)^2/2)/(15%)=0.341666667

d2=d1-volatility*sqrt(t)=0.341666667-0.15=0.191666667

N(d1)=0.63369912
N(d2)=0.575998341

C=S*N(d1)-Xe^(-rt)*N(d2)=20*0.63369912-20*e^(-4%)*0.575998341=1.605719938